Please answer this math problem. :)
When you reverse the digits in a certain two digit number you decrease its value by 27.
Find the number if the sum of its digits is 3?
Thank you so much!

Answer by mananth(16946)

When you reverse the digits in a certain two digit number you decrease its value by 27.
Find the number if the sum of its digits is 3?
...
let ten's digit be x
units digit be y
..
x+y = 3.................1
..
10y+x= 10x+y-27
9y-9x=-27
/9
y-x=-3...................2
add (1) & (2)
2y=-3+3
2y=0
y=0
So x = 3
The number is 30